# Life Sequence Transformer: Generative Modelling for Counterfactual Simulation

### v1

@@ Abstract
Social sciences rely on counterfactual analysis using surveys and administrative data, generally depending on strong assumptions or the existence of suitable control groups, to evaluate policy interventions and estimate causal effects. We propose a novel approach that leverages the Transformer architecture to simulate counterfactual life trajectories from large-scale administrative records. Our contributions are: the design of a novel encoding method that transforms longitudinal administrative data to sequences and the proposal of a generative model tailored to life sequences with overlapping events across life domains. We test our method using data from the Istituto Nazionale di Previdenza Sociale (INPS), showing that it enables the realistic and coherent generation of life trajectories. This framework offers a scalable alternative to classical counterfactual identification strategy—such as difference-in-differences and synthetic controls—particularly in contexts where these methods are infeasible or their assumptions unverifiable. We validate the model’s utility by comparing generated life trajectories against established findings from causal studies, demonstrating its potential to enrich labour market research and policy evaluation through individual-level simulations.
@@ 1 Introduction
Counterfactual inference is a cornerstone of scientific reasoning in disciplines such as political science Sekhon (2009), economics (Heckman, 2008), history (Levy, 2015), and psychology (Foster, 2010), and has also sparked ongoing methodological debate (Cartwright, 2007a). Counterfactual identification requires constructing hypothetical scenarios that resemble what would have happened in the absence of an event assumed to be causal (Rubin, 2005). Such a scenario allows quantifying the impacts of specific events—like treatments or policies—or the underlying causes of observed outcomes. Typical research questions include: What is the causal effect of increasing the minimum wage on household consumption (Alonso, 2022)? What are the drivers of civil conflict (Acemoglu et al., 2020)?
However, randomisation rarely occurs naturally and can be challenging to implement even in controlled settings, due to ethical constraints and substantial costs (Thomas, 2016). As a result, economists and other researchers have developed strategies and mathematical models to identify suitable control groups without relying on random treatment assignment. Despite being widely adopted, these observational methods often rely on specific assumptions about the data and the underlying causal mechanisms, which are typically difficult to verify in practice (for instance, see Appendix A).
Our contributions
This paper addresses a fundamental challenge in counterfactual inference: how to construct hypothetical scenarios that are both realistic and practically obtainable, especially when randomisation is unfeasible. We introduce a novel encoding method for life trajectories derived from administrative data in Section 3. Then, detail a Transformer foundational model for the generative task in Section 4. The model is trained on 1.6 million individual labour trajectories from INPS data, enabling scalable simulation of counterfactual life paths. In Section 5, our empirical analysis shows the generative potential of our encoding and architecture through two benchmarking exercises: evaluating its ability to simulate realistic life sequences aligned with traditional causal inference-driven findings in Section 5.2 and evaluating the coherence of generated out-of-sample trajectories in Section 5.3. Once trained, the model can generate realistic alternative life sequences by injecting or removing specific events from an individual’s observed trajectory, thus constructing counterfactuals without relying on external control groups.
@@ 2 Background
Our work builds on recent advances in Transformer-based sequence modelling for longitudinal labour and health data. In particularly, Life2Vec (Savcisens et al., 2024), CAREER (Vafa et al., 2024), and TransformEHR (Yang et al., 2023). Life2Vec showed the validity of encoder-only Transformer models for modelling socio-economic data and outperformed the predictive performance of standard models—like mortality tables, logistic regression, or recurrent neural networks—on individual’s future life outcomes (e.g., mortality) or time-invariant characteristics (e.g., personality nuances). However, their model lacks generative capability to simulate sequences of future tokens, and their data encoding cannot be easily adapted to a generative task. In contrast, our proposed encoding approach enables the training of a generative decoder, allowing for the simulation of counterfactual life sequences on a model architecture directly inspired by Life2Vec.
@@
The CAREER encoder-decoder Transformer model focused on job sequence generation with annual temporal resolution and no rich occupational attributes, such as work province or work title. We enhance this framework by introducing fine-grained temporal encoding and a richer set of characteristics that span labour and demographic dimensions. In CAREER, time is introduced with a timestep of one year, and there is no way of modelling the duration of events or the exact timing of occurrence. You could increase the timestep to monthly steps, but this would restrict generation to a fixed monthly grid, offering no control over when within the year events begin or how long they last. Our method of encoding time, by contrast, enables flexible temporal generation by allowing each life event to have a defined starting month and duration in months within the year.
@@
TransformEHR is an encoder-decoder Transformer generative model which encodes sequences of electronic health records and predicts the subsequent closest diagnoses (ICD-10 code). It includes demographic and time embeddings to model the numerical format of the visit date. TransformEHR was pre-trained to predict a patient’s future visits, complete ICD codes, given longitudinal ICD codes up to the current visit. Similarly to Life2Vec, the timing of each visit must be specified at inference time—meaning the model cannot decide when events occur, but only generate diagnoses for predefined dates. In contrast, our model allows the generation process to determine when events occur and the duration of each event. Crucially, our model supports variable event durations, which is essential for modelling life sequences that extend beyond instantaneous diagnoses.
Another relevant work is the REMI data encoding (Huang and Yang, 2020) within the Figaro encoder-decoder Transformer (von Rütte et al., 2023), which models sequences in the musical domain through structurally rich encodings. The encoding approach informs our design of temporally structured representations, which we adapt to model life events rather than notes. REMI encodes rhythmic structure using the beat-bar-phrase hierarchical structure in music into the sequence. Specifically, encoding a <Bar> event to indicate the beginning of a bar and the <Pos> events to point to certain locations within a bar. We use these representations in our encoding to represent the years and months when a given event occurs.
Finally, recent work on causal representation learning (e.g., Vafa et al. 2025, Gloeckle et al. 2024) explores debiased fine-tuning and interpretable interventions; however, it does not explore generative, individual-level counterfactuals. Our approach fills this gap by combining causal reasoning with sequence generation, enabling more nuanced and personalised estimations of treatment effects.
@@ 3 Data Encoding
We build on the work of Savcisens et al. (2024), where life events are concatenated into a sequence, with absolute and relative temporal information embedded as auxiliary attributes rather than forming the structural backbone of the sequence. Our encoding method imposes an explicit calendar-like structure on each life trajectory to support the sequential generation of future tokens. The entire observed work history is divided into yearly segments, each further subdivided into twelve monthly intervals, which are then populated with event information as needed. This structure mirrors the design used in music generation models (Huang and Yang, 2020): yearly segments correspond to musical bars, and monthly divisions set the tempo within each bar. Each life event acts like a musical note, characterised by attributes such as employment type, work intensity, sector, and qualification, much like musical notes vary in pitch, duration, and intensity.
For instance, if an individual remains employed at the same firm for 10 uninterrupted years, we encode this as 10 events—each representing one full year with a 12-month duration—and each is described using the same set of tokens. In musical terms, this resembles a sustained note repeated across multiple bars. Conversely, if a person holds a one-year contract but experiences a one-month interruption, the year is split into two separate events with different durations, but otherwise described by the same set of concept-tokens. When no life event occurs within a year, the structure still records an "empty" year, analogous to a silent bar in a piece of music, preserving the regular temporal rhythm of the sequence. Figure 1 illustrates the encoding pipeline. We consider each tabular record (i.e., each row in the administrative dataset) as an event or note and translate each column describing it into a concept-token. Then, we concatenate all the events for a given individual in chronological order and create an individual sequence.
@@
Figure 1: The step for converting data from Tabular format to the Generative model suitable format: (1) Convert relevant features to concept tokens (2) Look up relevant temporal information such as age and year (3) Concatenate token, time, and positional embeddings to the sequence.
@@
Each event includes embeddings regarding positional information: year or absolute position reporting the calendar year when the event occurred, age or relative position with the individual’s age at the time of the event, and the usual positional embedding used in Transformer models for the position of the token in the sequence. These embeddings represent an absolute timeline (calendar year), a relative perspective (age at the time of event), and the position of the token within the sequence itself. These three embeddings are not part of the generative task, instead, they are deterministic.
@@
Sequences start with the special token <BOL> (Beginning of Life). Individual events are then encoded using a fixed token order. <MONTH_X> marks the calendar month the event begins and helps separate multiple events within the same year. Next comes <TYPE_X> to indicate the event type (e.g., employment, unemployment), optionally followed by modifiers like <ATECO_X> (for sector) and <WRK_TITLE_X> (for job title), depending on the event. <DUR_X> defines the event’s duration in months and concludes the event. If more events follow in the same year, another <MONTH_X> appears; otherwise, the sequence ends with <EOY> (End of Year), or if no further events exist, the entire life sequence ends. Each sequence maintains a consistent ordering of these concept-tokens, ensuring a structured representation and facilitating the learning of this artificial language.
@@
While the experiments in this paper focus on labour-related life sequences, we emphasise that the proposed encoding scheme applies to other life domains, such as health and education. In the health domain, for instance, life sequences could represent diagnoses event attributes, annotated with the month of diagnosis and the duration of treatment or hospitalisation. Although a monthly resolution may appear too coarse for modelling certain health events, this limitation is mitigated by including intensity tokens <INTENSITY_X>, which capture the extent of engagement within each event’s duration. For a detailed description of the data and a labour life sequence example, see Appendix B.
@@
Time-invariant individual information—such as gender, place and date of birth—is also represented using concept-tokens (e.g., <F> or <M> for female and male). Background information can be placed at the beginning of the sequence—allowing the entire input to be modelled as a single stream of tokens—or process it separately through a separate encoding architecture. The latter design distinguishes between static demographic attributes and the temporal sequence of life events, which begins with the <BOL> token, and allows the temporal sequence to be more directly influenced by these central life attribute parameters.
4 Transformer Models for Life Sequences
This section describes the encoder–decoder Transformer architecture, illustrated in Figure 7. We are presenting an encoder–decoder and a decoder-only version of our model (see Figure 6) and will compare them in the experiments to highlight their respective advantages and limitations. The only architectural difference between them is that, in the decoder-only setup, the background information is appended to the input sequence rather than processed separately. Contrarily, in the encoder–decoder setup, time-invariant covariates are processed by the encoder and used by the decoder through cross-attention to generate life trajectories. Both models are trained autoregressively to predict the next event in the sequence, conditioned on both past events and background information.
@@ 4.1 Model Architecture
The decoder block follows a standard Transformer decoder design, consisting of three components: an embedding block, stacked Transformer-based decoder layers, and a task-specific output head. We make three key modifications to the standard architecture. First, we replace the conventional attention mechanism with Performer Attention (Choromanski et al., 2022), which reduces the number of trainable parameters and improves computational efficiency. Second, we incorporate Rotary Position Embeddings (RoPE) (Su et al., 2023), adapted to work with Performer Attention, to better capture temporal patterns in the sequence. Third, to stabilise training, we apply RMS Normalisation (Zhang and Sennrich, 2019) before the attention and feedforward layers. The decoder uses causal (or masked) self-attention, ensuring that each token can only attend to earlier positions in the sequence, and cross-attention allows it to attend to the output of the encoder—conditioning on background characteristics. During training, the model is optimised using a next-token prediction objective, learning to generate coherent event sequences conditioned on prior events and time-invariant variables.
The encoder block processes the time-invariant background information, which consists of only four concept-tokens. Given the short length of this input, we use standard full self-attention without approximation. Positional embeddings are omitted, as the tokens are unordered and independent. RMS Normalisation is applied before both the attention and feedforward layers. The encoder outputs a fixed-length representation that conditions the decoder via cross-attention, but we do not implement it on all decoders; in the other layers, only self-attention is used (like in Figure 6). We treat the number of layers with cross-attention as a tunable hyperparameter. Although background information is relevant, tuning showed that allocating parameters to more self-attention-only layers is a more efficient use of parameters.
@@
To encode temporal information—specifically, age and calendar year—we use the Time2Vec method (Kazemi et al., 2019), which allows the model to capture both linear and periodic aspects of time. This method uses trainable parameters $\omega$ and $\varphi$ to define a flexible sinusoidal function. We construct two separate time embeddings: one for age, denoted by $\mathcal{T}_{A}:\mathcal{A}\rightarrow\mathbb{R}^{d}$, and one for year, $\mathcal{T}_{Y}:\mathcal{Y}\rightarrow\mathbb{R}^{d}$. These embeddings are computed as:
$\mathcal{T}(x)[z]=\begin{cases}\tanh(w_{z}x+\varphi_{z})&\text{if }z\in[0,d//4)\text{ for Age or }z\in[0,d//2)\text{ for Year}\\ \sin(w_{z}x+\varphi_{z})&\text{if }z\in[d//4,d)\text{ for Age or }z\in[d//2,d)\text{ for Year}\end{cases}$ (1)
Initial iterations of the model used standard Time2Vec encoding, i.e. a linear component with trigonometric functions. However, we noticed that it led to poor out-of-sample performance: the model quickly lost coherence and failed to maintain the grammar of the life sequence when extrapolating beyond the training period. We hypothesised that this issue stemmed from the unbounded linear trend in the Time2Vec representation of age and year. Also, trigonometric components appeared to be mimicking linear trends—particularly for the year variable, but also for age.
@@ 4.2 Training Procedure
Our dataset consists of 1,652,877 individual life sequences. We divide them into training (70%), validation (15%), and test (15%) subsets. The models are trained for 15 to 20 epochs with early stopping. Each training batch contains 24 sequences, and gradients are accumulated over four steps before an optimiser step, resulting in an effective batch size of 96. This enables stable training without requiring large amounts of GPU memory.
@@
We sample hyperparameters using the Tree-Structured Parzen Estimator (Watanabe, 2023) and Optuna (Akiba et al., 2019). Experiments showed that larger models—especially with deeper networks—performed better. Without normalisation, the model misused the embedding space, often requiring unusually large embedding dimensions to compensate. With RMS normalisation, this effect was mitigated, although the model benefited from longer warmup phases compared to setups using only ReZero. Increasing depth proved more effective than scaling up feedforward or embedding dimensions, or the number of random features. For local attention, we set the window size to cover approximately three complete events, which empirically balanced context coverage and memory use. Ultimately, we selected hyperparameters that maximized model capacity while fitting within the 16GB of VRAM available on an RTX A4000 GPU: 10 decoder layers for decoder-only and 8 decoder layers for encoder-decoder, only 2 with cross-attention, with 8 attention heads, 6 local with window size of 36 and 2 global with projection of 256 random features; 1 encoder layer with 2 attention heads; embedding size of 240, and position-wise feed forward size of 960 on all layers. Both models have around 7.1 million tunable parameters.

### v2

# Life Sequence Transformer: Generative Modelling of Socio-Economic Trajectories from Administrative Data

@@ Abstract
Generative modelling with Transformer architectures can simulate complex sequential structures across various applications. We extend this line of work to the social sciences by introducing a Transformer-based generative model tailored to longitudinal socio-economic data. Our contributions are: (i) we design a novel encoding method that represents socio-economic life histories as sequences, including overlapping events across life domains; and (ii) we adapt generative modelling techniques to simulate plausible alternative life trajectories conditioned on past histories. Using large-scale data from the Italian social security administration (INPS), we show that the model can be trained at scale, reproduces realistic labour market patterns consistent with known causal relationships, and generates coherent hypothetical life paths. This work demonstrates the feasibility of generative modelling for socio-economic trajectories and opens new opportunities for policy-oriented research, with counterfactual generation as a particularly promising application.
@@ 1 INTRODUCTION
In fields such as physics (Geneva and Zabaras, 2022), chemistry (Kim et al., 2021), and health (Yang et al., 2023), researchers successfully use Transformer-based generative models to simulate surrogate dynamical systems under varying conditions and estimate event probabilities. In addition to their predictive capabilities, generative models have the potential to explore hypothetical system evolutions under different scenarios.
We extend this line of research to the social sciences by proposing a Transformer-based generative model for sequences of human life events. Longitudinal administrative records contain rich information on individuals’ trajectories across labour markets, health, education, and other socio-economic domains, yet these data are rarely treated as sequences for generative modelling. Through a novel encoding method of socio-economic events, we enable the training of a decoder-only Transformer model capable of generating plausible synthetic life trajectories.
Our model is trained on 1.6 million individual labour market histories from the Italian social security administration (INPS), described in Section 3. The representation of these histories as overlapping socio-economic events encoded into life sequences is introduced in Section 4. The architecture and training procedure of our life sequence Transformer are detailed in Section 5. The empirical analysis in Section 6 illustrates the generative potential of the framework through two benchmarking exercises: the accuracy of simulated sequences in capturing established causal nexus among life events (Section 6.2) and the predictive accuracy on out-of-sample trajectories (Section 6.3).
Once trained, the model can generate realistic continuations of partially input sequences and simulate alternative life paths. In other words, it can create counterfactuals, i.e., plausible “what if” scenarios that describe how events might have unfolded under different circumstances. This ability provides a scalable framework for constructing coherent counterfactual trajectories without the need for external control groups. In Section 7, we elaborate on how this approach can be integrated into experimental designs in the social sciences and outline promising directions for future research that build on our framework.
@@ 2 BACKGROUND
Our work builds on recent advances in Transformer-based sequence modelling for longitudinal labour and health data. In particularly, Life2Vec (Savcisens et al., 2024), CAREER (Vafa et al., 2024), and TransformEHR (Yang et al., 2023). Life2Vec tests the validity of encoder-only Transformer models for modelling socio-economic administrative data from Denmark. It outperforms the predictive performance of standard models, such as mortality tables, logistic regression, or recurrent neural networks, in predicting an individual’s future life outcomes (e.g., early mortality) or time-invariant characteristics (e.g., personality nuances). However, this model lacks generative capability to simulate sequences of future tokens, and the data encoding cannot be easily adapted to a generative task. In contrast, our proposed encoding approach enables the training of a generative decoder, allowing for the simulation of counterfactual life sequences on a model architecture directly inspired by Life2Vec.
@@
The CAREER encoder-decoder Transformer model focuses on job sequence generation with annual temporal resolution and no rich occupational attributes, such as firm size, social security benefits, or multiple job episodes within the same year. We enhance this framework by introducing fine-grained temporal encoding and a richer set of characteristics that span labour and demographic dimensions. In CAREER, time is introduced with a timestep of one year, and there is no way to model the duration of events or the exact timing of their occurrence. You could increase the timestep to monthly steps, but this would restrict generation to a fixed monthly grid, offering no control over when events begin or how long they last within the year. Our method of encoding time, by contrast, enables flexible temporal generation by allowing each life event to have a defined starting month and duration in months within the year.
@@
TransformEHR is an encoder-decoder Transformer generative model that encodes sequences of electronic health records and predicts the subsequent closest diagnoses (ICD-10 codes). It includes demographic and time embeddings to model the numerical format of the visit date. TransformEHR is pre-trained to predict a patient’s future visits with ICD codes, given longitudinal codes up to the current visit. Similarly to Life2Vec, the timing of each visit must be specified at inference time, meaning the model cannot decide when events occur, but only predict diagnoses for predefined dates. In contrast, our model allows the generation process to determine when events occur and their respective durations. Crucially, our model supports variable event durations, which is essential for modelling life sequences that extend beyond instantaneous diagnoses.
3 WHIP DATA
Our model is trained on the Work History Italian Panel (WHIP), a 1:15 sample of individuals drawn from the INPS database (Leombruni et al., 2010; Bena et al., 2012). INPS records cover most private-sector employees, quasi-dependent and self-employed workers, as well as periods of retirement and access to social benefits (e.g., unemployment or maternity leave). The WHIP provides detailed longitudinal employment and benefit histories for over 4 million individuals.
We use the WHIP database for the period 1990–2015 and apply sample selection criteria to obtain sequences that are sufficiently long to provide meaningful information. Specifically, we retain only individuals with complete information up to 2015, at least half of their potential labour force participation period observed, non-missing birth dates, at least five records in total, plausible entry ages (between 15 and 80), and a labour force attachment of at least 10%. This resulted in a sample of around 1.65 million life trajectories. A full description of the dataset and details of the selection procedure are provided in Appendix B.
Each labour event contains timing (start month and duration) and income, which is standardised into monthly real terms and discretised into 100 quantiles. Yearly income is the principal variable for the social security system, as it forms the basis for calculating contributions and benefits. Each income record is linked to a range of demographic, workplace, and labour-force-specific attributes. Demographic information includes sex, birth year, and month, and the macro-area of birth. Workplace characteristics cover the sector of activity, enterprise size, and workplace location at the provincial level.
Labour force information is recorded in terms of labour status (e.g., employee, self-employed, pensioner, unemployed) with additional modifiers depending on status. For employees, these include job title, work intensity, number of weeks of maternity leave, number of weeks of sick leave, and whether the contract is part-time or full-time. Together, these attributes enable a detailed reconstruction of individuals’ trajectories across work, unemployment, pensions, and benefit receipt. A full description of all 15 variables and encoding details is provided in Appendix B.1.
@@ 4 DATA ENCODING
Figure 1: The steps for converting data from Tabular format to the Generative model suitable format: (1) Convert relevant features to concept tokens (2) Look up relevant temporal information such as age and year (3) Concatenate token, time, and positional embeddings to the sequence.
We build on the work of Savcisens et al. (2024), where life events are concatenated in a sequential structure, with absolute and relative temporal information embedded as auxiliary attributes rather than forming the structural backbone of the sequence. Instead, our encoding method imposes an explicit calendar-like structure on each life trajectory to support the sequential generation of future tokens. The entire observed work history is divided into yearly segments, each further subdivided into twelve monthly intervals, which are then populated with event information as needed. Life events are represented as discrete units with attributes such as employment type, income quantiles, work intensity, sector, qualification, and duration, forming a sequence of socio-economic states.
For instance, if an individual remains employed at the same firm for 10 uninterrupted years, we encode this as 10 events, each lasting 12 months, described using the same set of descriptive tokens, and with a likely increase in income or a change in work title if promoted over the years. If a person holds a one-year contract but experiences a one-month interruption, the year is split into two separate events with different durations. Though the same set of concept-tokens is otherwise used to describe these events. When no event occurs within a year, the structure still records an “empty” year preserving the regular yearly pace of the sequence.
Figure 1 illustrates the encoding pipeline. We consider each tabular record (i.e., each row in the administrative dataset) as an event and translate each column describing it into a concept-token. Then, we concatenate all the events for a given individual in chronological order and create an individual sequence.
@@
Each event includes embeddings on positional information: year or absolute position reporting the calendar year when the event occurred, age or relative position with the individual’s age at the time of the event, and the usual positional embedding used in Transformer models for the position of the token in the sequence. Notice, these three embeddings are not part of the generative task; instead, they are deterministic.
@@
Sequences start with the special token <BOL> (Beginning of Life). Individual events are then encoded using a fixed token order. <MONTH_X> marks the calendar month the event begins and helps separate multiple events within the same year. Next comes <TYPE_X> to indicate the event type (e.g., employment, unemployment), optionally followed by modifiers like <ATECO_X> (for sector) and <WRK_TITLE_X> (for job title), depending on the event. <DUR_X> defines the event’s duration in months and concludes the block describing the event. If more events follow in the same year, another <MONTH_X> appears; otherwise, the <EOY> (End of Year) token concludes the description of the events in a year. If no further events ensue, the entire life sequence ends with an <EOL> (End of Life) token. Each sequence maintains a consistent ordering of these concept-tokens, ensuring a structured representation and facilitating the learning of this artificial language.
@@
While the experiments in this paper focus on labour-related life sequences, we emphasise that the proposed encoding scheme applies to other life domains, such as health and education. In the health domain, for instance, life sequences could represent concatenations of diagnostic attributes, annotated with the month of diagnosis and the duration of treatment or hospitalisation. Although a monthly resolution may appear too coarse for modelling certain health events, this limitation is mitigated by including intensity tokens <INTENSITY_X>, which capture the extent of engagement within each event’s duration. For a detailed description of the data and a labour life sequence example, see Appendix B.
@@
Background time-invariant information, such as gender, place, and date of birth, is also represented using concept-tokens (e.g., <F> or <M> for female and male). This information is placed at the beginning of the sequence, forming the entire input as a single stream of tokens. The encoding distinguishes between static demographic attributes and the temporal sequence of life events, which begins with the <BOL> token.
5 TRANSFORMER MODEL
We propose a decoder-only Transformer architecture, see Figure 7, trained autoregressively to predict the next token in the sequence, conditioned on past events and background information.
@@ 5.1 Model Architecture
The model follows a Transformer decoder-only design, consisting of three components: an embedding block, stacked Transformer-based decoder layers, and a task-specific output head. We make three key modifications to the standard architecture. First, we replace the conventional attention mechanism with Performer Attention (Choromanski et al., 2022), which reduces the number of trainable parameters and improves computational efficiency. Second, we incorporate Rotary Position Embeddings (RoPE) (Su et al., 2023), adapted to work with Performer Attention, to better capture temporal patterns in the sequence. Third, to stabilise training, we apply RMS Normalisation (Zhang and Sennrich, 2019) before the attention and feedforward layers. The decoder uses causal (or masked) self-attention, ensuring that each token can only attend to earlier positions in the sequence. During training, the model is optimised using a next-token prediction objective, learning to generate coherent event sequences conditioned on prior events and time-invariant variables.
@@
To encode temporal information, i.e., age and calendar year, we use the Time2Vec method (Kazemi et al., 2019), which allows the model to capture both linear and periodic aspects of time. This method uses trainable parameters $\omega$ and $\varphi$ to define a flexible sinusoidal function. We construct two separate time embeddings: one for age, denoted by $\mathcal{T}_{A}:\mathcal{A}\rightarrow\mathbb{R}^{d}$ with $Z:=4$, and one for year, $\mathcal{T}_{Y}:\mathcal{Y}\rightarrow\mathbb{R}^{d}$ with $Z:=2$. These embeddings are computed as:
$\mathcal{T}(x)[z]=\begin{cases}\tanh(w_{z}x+\varphi_{z})&\text{if }z\in[0,d//Z)\\ \sin(w_{z}x+\varphi_{z})&\text{if }z\in[d//Z,d)\end{cases}$ (1)
Initial iterations of the model used standard Time2Vec encoding, i.e., a linear component with trigonometric functions. However, we noticed that it led to poor out-of-sample performance: the model quickly lost coherence and failed to maintain the grammar of the life sequence when extrapolating beyond the training period. We hypothesised that this issue stemmed from the unbounded linear trend in the Time2Vec representation of age and year. Additionally, trigonometric components appeared to mimic linear trends, particularly for the year variable, and also for age.
@@ 5.2 Training Procedure
Table 1: Test set metrics for next-token prediction, macro-averaged across the vocabulary, conditioning on 0, 5, 1 and 10 initial years of sequence.
Accuracy Precision Recall F1 Score Perplexity
0 known years 0.6225 0.6868 0.6225 0.6427 0.7288
1 known year 0.6773 0.7216 0.6773 0.6895 0.6937
5 known years 0.6883 0.7285 0.6883 0.6991 0.6727
10 known years 0.7039 0.7409 0.7039 0.713 0.6466
Our dataset consists of 1,652,877 individual life sequences. We divide them into training (70%), validation (15%), and test (15%) subsets. The models are trained for 15 to 20 epochs with early stopping. Each training batch contains 18 sequences, and gradients are accumulated over five steps before an optimiser step, resulting in an effective batch size of 90.
@@
We sample hyperparameters using the Tree-Structured Parzen Estimator (Watanabe, 2023) and Optuna (Akiba et al., 2019). Experiments showed that deeper models performed better. Without normalisation, the model misuses the embedding space, often requiring unusually large embedding dimensions to compensate. With RMS normalisation, this effect was mitigated, although the model benefited from longer warmup phases compared to setups using only ReZero. Increasing depth proved more effective than scaling up feedforward or embedding dimensions, or the number of random features. For local attention, we set the window size to cover approximately three complete events, which empirically balanced context coverage and memory use. Ultimately, we selected hyperparameters that maximized model capacity while fitting within the 16GB of VRAM available on an RTX A4000 GPU: 10 decoder layers, with 8 attention heads, 6 local with a window size of 36 and 2 global with projection of 256 random features; embedding size of 240, and position-wise feed forward size of 960 on all layers. The model has around 10.1 million tunable parameters.